The following table contains information on matched sample values whose differences are normally distributed. (You may find it useful to reference the appropriate table: z table or t table)

Number Sample 1 Sample 2
1 17 20
2 12 12
3 21 22
4 21 20
5 16 21
6 14 16
7 17 18
8 17 20


a.
Construct the 90% confidence interval for the mean difference μD. (Negative values should be indicated by a minus sign. Round intermediate calculations to at least 4 decimal places and final answers to 2 decimal places.)
Answer #1

The following table is obtained:

Sample 1 Sample 2 Difference = Sample 1 - Sample 2
17 20 -3
12 12 0
21 22 -1
21 20 1
16 21 -5
14 16 -2
17 18 -1
17 20 -3
Average 16.875 18.625 -1.75
St. Dev. 3.091 3.249 1.909
n 8 8 8

and the sample size is n = 8. For the score differences, we have

D = -1.75sp= 1.909

We need to construct the 90% confidence interval for the population mean μ. The following information is provided:

Sample Mean -1.75
Sample Standard Deviation 1.909
Sample Size 8

The critical value for α = 0.1 and df = n-1 = 7 degrees of freedom is te = 21-a/2;n-1 = 1.895. The corresponding confidence interval is computed as shown below:

ca (F-A*、A+tx =)

CI = (-1.75 -1.895– 1.909 1.909 -1.75 +1.895– V8 V8 : 1.75

CI = (-3.029, -0.471)

Therefore, based on the data provided, the 90% confidence interval for the population mean is -3.029 < μ < −0.471, which indicates that we are 90% confident that the true population mean μ is contained by the interval (-3.029, -0.471)